Protein X is synthesized in vitro using 35S-methionine that has a specific radioactivity of 25 Ci/mmole. Protein X contains 6 methionine residues. After synthesis, 10.0 microliters of the solution is removed, and the protein is separated from the unincorporated (free) amino acids. This protein sample is found to contain 3.5 x 104 CPM in an instrument that has a counting efficiency of 35%. What is the concentration of protein X in this sample in units of molarity.
Answer #1

Protein X using 35S-methionine.

Specific Radioactivity of 35-S = 25 Ci/mmole

1 Ci = 2.22*106 disintegration per minutes

So, Specific Radioactivity of 35-S = 25*2.22*106 disintegration per minutes / mmole

= 5.55 *106 disintegration per minutes / mmole

Specific Radioactivity of 35-S      =  5.55 *109 disintegration per minutes / mole (1 mmol = 10-3 mol )

This protein sample had = 3.5 x 104 CPM (count per minute)

Counting efficiency of Instrument = 35 % (It is ratio between the number of particles/photons counted with a counter and the number of particles/ photons of the same type and energy emitted by the source)

count per minute =  disintegration per minutes

So,   disintegration per minutes in protein sample = 105  disintegration per minute.

Sample volume = 10.0 Microliters = 10-5 L.

or disintegration per minutes in protein sample =  1010  disintegration per minute / L.

Since, Protein X contains 6 Methionine Residue.

disintegration per methionine residue = 1.67 * 109  disintegration per minute / L.

Concentration of protein X in this sample =

(1.67 * 109  disintegration per minute / L)/ 5.55 *109 disintegration per minutes / mole

= 0.30 mole / L = 0.30 M